OpenStudy (anonymous):

how do I simplify:

OpenStudy (anonymous):

\[\log_{ 3 } (1/3)\]

OpenStudy (anonymous):

beneath the log is a 3!

OpenStudy (anonymous):

I have done logs before. log 3 (1/3) = x (what you are trying to find out) 3^x = 1/3 x = ?

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

you can do it like this: log3(1/3)=log3(3^(-1))=-1(log3(3))=-1(1)=-1